2 protons are located 6.0 x 10^-15 m away from one another. What is the electrostatic potential energy of the pair?
Answer #1

280 9= +16*109€, 7 = 6*10-15m u= K2² = 9x107x (1.6x10-19)² (6x1015) 13). 4 = 3-84210-19 Joule = 3.84x10-19 Joule